A sales clerk receives a monthly salary of $⁢700 plus a commission of 5% on all sales over $⁢4000. What did the clerk earn the month that he sold $⁢11,800 in merchandise?
Answer by stanbon(75887)

A sales clerk receives a monthly salary of $⁢700 plus a commission of 5% on all sales over $⁢4000. What did the clerk earn the month that he sold $⁢11,800 in merchandis
-------------------------
Earning = 700 + 0.05(11800-4000) = 700 + 0.05*7800 = $1090.00
----------
Cheers,
Stan H.